Three fair dice are shaken and rolled onto a table. What is the probability that all three dice will display an odd number of spots on their top faces?
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

Three fair dice are shaken and rolled onto a table. What is the probability that all three dice will display an odd number of spots on their top faces?
-----
Total # of elements in sample space: 6^3 = 216
# of ways to succeed
1,3,5::: 6 ways
----
P(3 odd faces) = 6/216 = 1/36
===============================
Cheers,
Stan H.
===========

Answer by Edwin McCravy(20054)

Three fair dice are shaken and rolled onto a table. What is the probability that all three dice will display an odd number of spots on their top faces?
The other tutor's answer is wrong. The answer is 1/8.  
Because each of the following 8 cases is equally likely
because a die has 3 even spots and 3 odd spots.  The
8th one below is the 1 successful one among the 8.

    die1  die2  die3 
--------------------
1.  EVEN  EVEN  EVEN
2.  EVEN  EVEN  ODD
3.  EVEN  ODD   EVEN
4.  EVEN  ODD   ODD
5.  ODD   EVEN  EVEN 
6.  ODD   EVEN  ODD
7.  ODD   ODD   EVEN
8.  ODD   ODD   ODD

Edwin